OpenStudy (anonymous):

Multiply and simplify:

OpenStudy (anonymous):

\[\frac{ x^2-81 }{ x^2+18+81 } * \frac{ x+9 }{ x-9 }\]

OpenStudy (anonymous):

answers: a) x+9 ---- x-9 b) 1 --- x+9 c) 0 d) 1

OpenStudy (***[isuru]***):

hello u can write the right hand side numerator using the "difference of 2 squares" \[x ^{2} - 81 = (x -9)(x + 9)\] and u can factorize the denominator of the right hand side as \[x ^{2} + 18x + 9 = (x + 9 )^{2}\] now re-write the expression by substituting those values and u could simplify it easily Did u get it ?

OpenStudy (tkhunny):

Have you considered factoring everything?

OpenStudy (yttrium):

Just factor quadratics, and do some cancellation and see what will happen. :)